Prove that lim (Xn)=0 if and only if lim(|Xn|)= 0. Give an example to show that the convergence of (|Xn|)= 0 need not imply the convergence of (Xn)= 0.
1
Expert's answer
2019-08-26T10:35:16-0400

"\lim\limits_{n\to\infty} X_n=0 \leftrightarrow \forall \varepsilon>0 \ \exists N \ \forall n>N \ |X_n-0|<\varepsilon."

"\lim\limits_{n\to\infty} |X_n|=0 \leftrightarrow \forall \varepsilon>0 \ \exists N \ \forall n>N \ \bigl||X_n|-0\bigr|<\varepsilon."

Since "|X_n-0|=\bigl||X_n|-0\bigr|" , we have "\lim\limits_{n\to\infty} X_n=0 \Leftrightarrow \lim\limits_{n\to\infty} |X_n|=0" .

Consider sequence "X_n=(-1)^n." It is divergent one, but sequence "|X_n|=1" is convergent.